OpenStudy (anonymous):

Find the center of this hyperbola: -4x^2+y^2-8x-12y+36=0

OpenStudy (anonymous):

ya, but this equasion seems to break half-way through

OpenStudy (nottim):

arghhh don't complex me with the simplicities of maths!

OpenStudy (anonymous):

I try my best.

OpenStudy (anonymous):

honestly though, can someone work through this with me?

OpenStudy (nottim):

bahhahahaha

OpenStudy (anonymous):

go away. lol

OpenStudy (anonymous):

stoppit

OpenStudy (anonymous):

dont you freaken post that....just backspace now

OpenStudy (nottim):

no more?

OpenStudy (anonymous):

more please actually

OpenStudy (nottim):

I'm all out.

OpenStudy (anonymous):

then no medal for you...

OpenStudy (anonymous):

ZZ, i need help on like 4 hyperbola Q's, cause i can do the ellipses

OpenStudy (anonymous):

LOL! "Do you mean a parabola? the center of hyperbola is the letter r. " thats what chacha said

OpenStudy (nottim):

EEK

OpenStudy (zzr0ck3r):

sec

OpenStudy (zzr0ck3r):

-4x^2+y^2-8x+36 = 0 -4x^2-8x-4+y^2-12y+36+4=0 -4(x^2+2x+1) + (y-6)^2 = -4 (x+1)^2^2+(y-6)^2 = 1

OpenStudy (zzr0ck3r):

got the rest?

OpenStudy (nottim):

Yayy. Real answer.

OpenStudy (anonymous):

so its (x+1)^2+(y-6)^2=1 How do i get the center ? -1,6?

OpenStudy (zzr0ck3r):

yep

OpenStudy (anonymous):

ok, so how do i get vertices? now in standard form: (x+1)^2/1 - (y-6)^2/4=1

OpenStudy (zzr0ck3r):

where did th e/4 come from?

OpenStudy (anonymous):

thats what it gives me though. (x+1)^2/1 - (y-6)^2/4 how do i find the vertices?

OpenStudy (zzr0ck3r):

o yeah sorry I forgot something -4x^2+y^2-8x+36 = 0 -4x^2-8x-4+y^2-12y+36+4=0 -4(x^2+2x+1) + (y-6)^2 = -4 (x+1)^2+(y-6)^2/-4 = (x+1)^2-(y-6)^2/4=1

OpenStudy (zzr0ck3r):

I forget vertices, one sec

OpenStudy (anonymous):

no no its cool, its kinda hard without seeing it

OpenStudy (anonymous):

i need help finding the Foci for that problem